OpenStudy (anonymous):

Factor: x2 + 11x + 28 A. (x + 7)(x + 4) B. (x + 7)(x - 4) C. (x + 8)(x + 3) D. (x + 14)(x + 2)

OpenStudy (anonymous):

@Aureyliant

OpenStudy (anonymous):

Sure.. give me two mins

OpenStudy (anonymous):

okay :) @Aureyliant

OpenStudy (anonymous):

Okay, so first off, I need to know your understanding of factoring trinomials before I give you the straight out answer. (I give explanations too but I just need to know you understand this, too)

OpenStudy (anonymous):

okay i didnt get it that much @Aureyliant

OpenStudy (anonymous):

Trying to factor by splitting the middle term 2.1 Factoring x2+11x+28 The first term is, x2 its coefficient is 1 . The middle term is, +11x its coefficient is 11 . The last term, "the constant", is +28 Step-1 : Multiply the coefficient of the first term by the constant 1 • 28 = 28 Step-2 : Find two factors of 28 whose sum equals the coefficient of the middle term, which is 11 . -28 + -1 = -29 -14 + -2 = -16 -7 + -4 = -11 -4 + -7 = -11 -2 + -14 = -16 -1 + -28 = -29 1 + 28 = 29 2 + 14 = 16 4 + 7 = 11 That's it Step-3 : Rewrite the polynomial splitting the middle term using the two factors found in step 2 above, 4 and 7 x2 + 4x + 7x + 28 Step-4 : Add up the first 2 terms, pulling out like factors : x • (x+4) Add up the last 2 terms, pulling out common factors : 7 • (x+4) Step-5 : Add up the four terms of step 4 : (x+7) • (x+4) Which is the desired factorization Final result : (x + 7) • (x + 4)

OpenStudy (anonymous):

Read this and ask me any Q's or tell me anywhere you're confused.

OpenStudy (anonymous):

wait did you mean 2x or x+2

OpenStudy (anonymous):

x2 is aka x^2

OpenStudy (anonymous):

i get it now @Aureyliant

OpenStudy (anonymous):

AWESOME!! I love "hearing" that from people I help.

OpenStudy (anonymous):

Need anymore help? Just tag me :)